# Silhouette Aggregation: From Micro to Macro

### v2

@@ Abstract
Silhouette coefficient is an established internal clustering evaluation measure that produces a score per data point, assessing the quality of its clustering assignment. To assess the quality of the clustering of the whole dataset, the scores of all the points in the dataset are either (micro) averaged into a single value or averaged at the cluster level and then (macro) averaged. As we illustrate in this work, by using a synthetic example, the micro-averaging strategy is sensitive both to cluster imbalance and outliers (background noise) while macro-averaging is far more robust to both. Furthermore, the latter allows cluster-balanced sampling which yields robust computation of the silhouette score. By conducting an experimental study on eight real-world datasets, estimating the ground truth number of clusters, we show that both coefficients, micro and macro, should be considered.
@@ 1 Introduction
The silhouette coefficient [1] serves as a widely used measure for assessing the quality of clustering assignments of individual data points. It produces scores on a scale from $-1$ to $1$ reflecting poor to excellent assignments, respectively. In real world applications, where it is widely accepted [2, 3, 4], it is common practice to average these scores to derive a single (micro-averaged) value for the entire dataset. However, this averaging approach can produce misleading results in certain scenarios, such as cluster imbalance or the presence of data points lying between or far from the clusters (background noise).
In this study, we advocate averaging the silhouette values both across clusters (macro-averaging) and across data points (micro-averaging) to obtain a single score for evaluating the clustering assignments of the entire dataset. The micro-averaged silhouette index is the originally proposed aggregation strategy [1] and the default implementation in scikit-learn, which have made it a very popular approach, e.g., to define the number of clusters. In this work, we show that by assigning equal weight to each cluster, the former is more effective in imbalanced or noisy cases.
To substantiate our claim, we have conducted a series of experiments on synthetic imbalanced and noisy clustering problems, and by estimating the number of clusters in eight real-world datasets. Our results show that the macro-averaged silhouette coefficient is more robust concerning noise, imbalance, and sample size.
@@ 2 The Silhouette Coefficient
Assessing the quality of a clustering solution ideally requires human expertise [8]. However, finding human evaluators could be hard, expensive and time-consuming (or even impossible for very large datasets). An alternative approach is to use clustering evaluation measures, which can be either external (supervised) or internal (unsupervised)  [9]. The former, as the name suggests, use external information (e.g., classification labels) as the ground truth cluster labels. Well known external evaluation measures are Normalised Mutual Information (NMI) [10], Adjusted Mutual Information (AMI) [11], Adjusted Rand Index (ARI) [12, 13], etc. External information, however, is not typically available in real-world scenarios. In such cases we resort to internal evaluation measures, which are solely based on information intrinsic to the data. Although other internal evaluation measures have been proposed [14, 15], we focus on the most commonly-employed, and successful one based on an extensive comparative study [16], which is the silhouette coefficient [1].
@@
Figure 1: Illustration of the elements involved in computing the silhouette score $s(x_{i})$.
@@
3 Aggregating Silhouette
Since the silhouette coefficient provides a score that grades the cluster assignment of a data point, in order to obtain a single score for the whole partition of dataset $X$, the individual silhouette scores have so far either (micro) averaged at the data point level or macro-averaged at the cluster level.
3.1 The Micro-averaged Silhouette Index
Micro-averaging silhouette at the point level is defined as follows:
@@ 3.1 The Micro-averaged Silhouette Index
This is the proposed averaging strategy by the study that introduced silhouette [1] and it is adopted by the widely-used library of scikit-learn, making it the typical approach employed in literature [17]. However, as we show in this work, this aggregation strategy has two weaknesses. First, it is not effective in the case of imbalanced clusters. Second, it is sensitive to the presence of background noise, i.e., data points lying in the space between the clusters.
3.2 The Macro-averaged Silhouette Index
(a) Synthetic dataset.
(b) Optimal clustering.
(c) Suboptimal clustering.
Figure 2: (a) A synthetic dataset, (b) its optimal clustering solution, (c) suboptimal clustering solution. The micro-averaged silhouette score $S_{micro}$ of the optimal solution is 0.74, which drops to -0.20 when the points in all clusters, except for the compact (low-variance) cluster, in the center are randomly assigned a cluster label.
When clusters are perfectly balanced, the sample mean (i.e., micro-averaging) is a reasonable aggregation strategy. This assumption, however, of clusters being perfectly balanced, cannot be guaranteed in the real world, where clusters are often imbalanced (see, for example, Table 1). In this case, micro-averaging is not effective, as will be discussed next in more detail. By contrast, the cluster-level (macro) aggregation strategy is more robust. In this strategy, a silhouette score $S_{c}$ is computed for each cluster $C_{i}$ as follows [^1] We note that the R implementation of the silhouette score employs macro-averaging. To the best of our knowledge, very few studies in literature deviate from the typical approach and follow this paradigm [18].:
@@ 3.2 The Macro-averaged Silhouette Index
Hence, for each cluster $C_{i}$ a single silhouette value $S_{c_{i}}$ is computed, known as cluster mean silhouette. This value can be interpreted as a score that measures how compact and well separated a cluster given a clustering solution. Consequently, we end up with a set of $K$ silhouette values {$S_{C_{1}},\ldots,S_{C_{K}}$}, where $K$ is the number of clusters in the solution. To evaluate the clustering solution as a whole, the mean aggregates these $K$ scores into a single value. Although the mean is a straightforward option, we note that other statistics can also be used for this aggregation (e.g., min, max, median). More formally, this macro-averaged silhouette variant (i.e., equally weighing the clusters) is defined as follows:
@@
3.3 An imbalanced clustering example
Consider the synthetic dataset shown in Figure 1(a), which consists of $12$ Gaussian clusters with $100$ data points each. The clusters are of different variance, ranging from low to high values. Figure 1(b) presents the perfect clustering, where points in the same cluster are coloured similarly. The micro-averaged silhouette score $S_{micro}$ of this perfect clustering is $0.74$. Figure 1(c) shows the same data points as in Figure 1(b), but a random cluster assignment is hypothesised for all data points, except for those belonging to the compact low-variance cluster (nucleus) lying at the center of the dataset. For this suboptimal clustering, $S_{micro}$ drops to $-0.20$. Then, we disturb the cluster balance by adding points to the nucleus cluster, using a very small variance to preserve its dense structure, and we evaluated how the increase in cluster imbalance affects the $S_{micro}$ score. Figure 3 shows that the score increases monotonically with the nucleus size and it even exceeds the score of the correct partition for size above $5,000$ data points. $S_{macro}$, on the other hand, has a value of $-0.20$ at all cases of this random cluster assignment and a value of 0.74 for the perfect clustering.
Figure 3: Micro- and macro-averaged silhouette (vertically) as a function of the size of the nucleus cluster (Figure 1(c)). The score for the optimal clustering (Figure 1(b)) is shown in yellow. In black is shown the highest possible silhouette score.
3.4 Cluster-balanced sampling for silhouette computation
@@
The computation of the silhouette coefficient for all the $N$ points in a dataset requires the computation of a pairwise distance matrix at the cost of $\mathcal{O}(N^{2})$ operations. This is demanding in terms of computational and space complexity and, hence, not scalable for large datasets [19]. The typical approach to tackle this problem is to compute the silhouette score using a uniformly selected subsample of the dataset. In a cluster-imbalanced problem, the typical (uniform across data points) sampling may favour the major cluster and may even disregard completely one of the minor clusters. For example, considering the dataset of Figure 2 that contains the very dense nucleus cluster, the silhouette score was often impossible to compute when using the typical sampling strategy, since only data points from the nucleus cluster were sampled.
@@
To solve this problem, following the intuition of the macro-averaged silhouette score, we propose creating a subsample of size $L$ for computing the macro-averaged silhouette score by uniformly selecting a subset of $L/K$ points from each cluster $C_{i}$, where $i=1,\ldots,K$. In this way, we ensure that all the clusters contribute a sufficient number of data points to the subsample, and a robust macro-averaged silhouette score computation.
Figure 4: Macro-averaged silhouette values (vertically) as a function of the size of the typical uniform across data points (red) and the proposed balanced per-cluster (green) sample. The median (solid line) and the inter-whisker range (shadowed) from $30$ Monte Carlo simulations are shown.
Figure 4 presents a comparison between the typical (i.e., uniformly across data points) and the proposed cluster-balanced sampling of the macro-averaged silhouette score. The median for both strategies converges to the value achieved when using the whole dataset (yellow, dashed line), as expected. However, when using the typical uniform sampling strategy (red line), the macro-averaged silhouette score considerably varies (red shadow), especially for small sample sizes. By contrast, the variation is small when using the balanced strategy (green shadow), converging faster to a small value as the sample size increases.
3.5 Estimating the true number of clusters
One of the typical uses of silhouette is to estimate the number of clusters. This is achieved by solving the clustering problem for several $K$ values and selecting the solution with the maximum silhouette score [20].
To study the effect of imbalance on estimating the number of clusters we consider again the synthetic dataset of Figure 2. Each cluster consists of $100$ data points, except for the compact nucleus cluster that contains $10,000$ points. Using this dataset and for number of clusters $k\in[2,30]$, we have applied the global $k$-means++ [21], which is a fast variant of the global $k$-means method [22] employing the $k$-means++ [23] seed selection.
@@ 3.5 Estimating the true number of clusters
Figure 5: Estimating of number of clusters $k$ using the micro-averaged and macro-averaged silhouette using a synthetic dataset with imbalanced clusters.
Figure 5 shows the silhouette score of the solutions for varying $k$, when it is macro- (green line) and micro-averaged (red line). A star indicates the maximum per line. The macro-averaged score always finds the true number of clusters (yellow dashed vertical line), despite the heavy cluster imbalance. In contrast, the default micro-averaged scores are saturated near $0.95$ and a plateau for $k\in[6,15]$. The lack of a clear peak leads to the wrong estimation for the true number of clusters.
3.6 Sensitivity to background noise
Besides its sensitivity on cluster imbalance, shown in the previous paragraphs, we present another weakness of the micro-averaged silhouette score that is related to the existence of background noise, i.e., points lie between clusters or far from the clusters.
Figure 5(a) depicts four well-separated Gaussian clusters, each comprising $200$ data points. We introduced background random noise to this dataset, by sampling $n$ points from a uniform distribution $\epsilon_{n}\sim\mathcal{U}(a,b)$, with $a,b$ chosen so that the samples are spread uniformly in the data space, as shown in Figure 5(d). If $n$ noisy points are added, we define the relative level of background noise that is added to a dataset $X$ as $\frac{n}{n+|X|}\times 100\%$, As an example, Figure 5(b) (5(c)) presents the original dataset with 25% (50%) of background noise.
We generated several datasets (i.e., versions of the original dataset) by increasing the noise level. To each such dataset we applied global $k$-means++, with $k$ ranging from 1 to 30, and we computed the micro- and macro-averaged silhouette scores per $k$. The $k$ with the highest score was considered as the estimated correct number of clusters per dataset. Figure 5(d) presents the number of clusters estimation with respect to the noise level, where it can be seen that the macro-averaged score always provides the correct number of clusters. In contrast, the estimations using the micro score are severely affected as the noise level increases.
(a) Original dataset
(b) $25\%$ is background noise
(c) $50\%$ is background noise
(d) Estimated number of clusters (vertically) per background noise level (horizontally) using micro- and macro-averaged silhouette scores
Figure 6: Synthetic dataset with four Gaussian clusters (a), focalised for visibility. Random background noise is added, consisting of 25% (b) and 50% (c) of the data while the estimated number of clusters is depicted per background noise level included in the data (d).
(a) pendigits
(b) olivetti faces
(c) glass

### v3

# Revisiting Silhouette Aggregation

@@ Abstract
Silhouette coefficient is an established internal clustering evaluation measure that produces a score per data point, assessing the quality of its clustering assignment. To assess the quality of the clustering of the whole dataset, the scores of all the points in the dataset are typically (micro) averaged into a single value. An alternative path, however, that is rarely employed, is to average first at the cluster level and then (macro) average across clusters. As we illustrate in this work with a synthetic example, the typical micro-averaging strategy is sensitive to cluster imbalance while the overlooked macro-averaging strategy is far more robust. By investigating macro-Silhouette further, we find that uniform sub-sampling, the only available strategy in existing libraries, harms the measure’s robustness against imbalance. We address this issue by proposing a per-cluster sampling method. An experimental study on eight real-world datasets is then used to analyse both coefficients in two clustering tasks.
@@ 1 Introduction
The silhouette coefficient [1] serves as a widely used measure for assessing the quality of clustering assignments of individual data points. It produces scores on a scale from $-1$ to $1$ reflecting poor to excellent assignments, respectively. In real world applications, where it is widely accepted [2, 3, 4], it is common practice to average these scores to derive a single (micro-averaged) value for the entire dataset. This is the originally proposed aggregation strategy [1] and the only implementation in the popular scikit-learn machine learning library in Python. An alternative aggregation strategy, however, is to average the Silhouette scores per cluster and then (macro) average across clusters, but our exploration of the related literature shows that this is a strategy that is rarely used in published studies. This is an alarming finding, because micro-averaging, e.g., in a classification context, is known to be sensitive to class imbalance [5, 6].
In this work, we focus on the effect of micro-averaging to a very well known internal cluster validity index, addressing the following research question RQ1: Is micro-averaging, which is the typical strategy to aggregate Silhouette scores in cluster analysis, sensitive to cluster imbalance? We answer this question using synthetic data, showing that micro-averaging Silhouette can produce misleading results for clustering solutions with imbalanced clusters. By contrast, we show that macro-averaging, which is rarely used in literature, is considerably more robust to this issue, because it assigns equal weight per cluster while disregarding its size.
In cluster analysis, Silhouette scores are often subsampled before being aggregated to yield a single score for a large dataset. This is a particularly useful step for computationally expensive tasks, as for example when assessing clustering solutions for a varying number of clusters to select the optimal [7]. By evaluating existing libraries in Python and R, we observe that only uniform sampling is implemented, which makes us focus on a second research question RQ2: Is uniform sampling suitable when macro-averaging or is its robustness against cluster-imbalance put at stake? The answer is the latter. Theoretically, in an extremely imbalanced dataset, the smallest cluster could even disappear when sub-sampling uniformly, which would exclude one (equal) factor from the macro-average. We address this issue, by proposing a novel per-cluster sampling strategy, which we show that it best suits macro-averaging of Silhouette scores.
Overall, the contributions of this study can be summarised in the following three aspects:
  • We compare two aggregation strategies that can be used to compute a Silhouette score for a dataset, showing that the typical micro-averaging strategy is problematic for imbalanced datasets.
  • We introduce a per-cluster sampling strategy, which should be the one used along with macro-averaging.
  • We quantify the sensitivity of micro-averaged Silhouette on imbalanced synthetic data, and we analyse two real-world imbalanced datasets on which the macro average should be preferred.
The remainder of this study comprises the related work (§2), a description of the Silhouette Coefficient (§3) and the aggregating strategies (§4), followed by an investigation on synthetic data (§5.1), an experimental study on real-word data (§5), and closing with remarks and future directions. Our code is publicly available in https://github.com/ipavlopoulos/revisiting-silhouette-aggregation.
2 Related Work
The typical approach
The vast majority of published studies employ micro-averaging to report the Silhouette Coefficient. In [8, 9, 10], the authors focus on the number of clusters estimation problem. In [11], they use the term Average Silhouette Width to refer to the micro-averaged Silhouette score. The authors of [12] explicitly report the implementation of scikit-learn that employs the micro-averaging strategy. The authors of [13], proposed a clustering algorithm that divides recursively the clustered dataset based on the maximization of the silhouette index. Similarly with the rest studies, they used the micro-average strategy, which they called as summation of silhouettes.
Exceptions to the rule
In [14], the author observes that SPSS and R employ different implementations of the Silhouette score. They note that the latter is using micro-averaging and is the correct out of the two. We observe, however, that other libraries (packages) do not necessarily follow this paradigm. ClusterCrit, [^1] https://cran.r-project.org/web/packages/clusterCrit/vignettes/clusterCrit.pdf for example, compute cluster mean silhouette scores that they average to yield the final index, but this is in fact a macro-averaged score. Without any study in published literature to assess the two strategies, we argue that this is considerably problematic, because, as we show (§5), results reported with the two strategies on the same data may not be comparable. Notably, we could only detect just one study using (and explicitly stating) the macro-averaging implementation [15].
Filling the gap
Micro-averaging is the typical and widely-used approach when aggregating Silhouette, with macro-averaging being considerably overlooked in literature of cluster analysis. Absent in existing literature is also a comparative study between the two aggregation strategies, a gap that is being bridged for classification tasks [5, 6]. Our study of existing macro-averaging implementations (ClusterCrit) reveals that only uniform sampling is employed, often used for the application on large datasets. We observe, however, that uniform sampling cancels the benefits of macro-averaging (i.e., in cluster-imbalanced spaces), because the measure reduces effectively to micro-averaging. We address this gap by proposing an alternative sampling approach, well-suited to macro-averaging.
@@ 3 The Silhouette Coefficient
Assessing the quality of a clustering solution ideally requires human expertise [19]. However, finding human evaluators could be hard, expensive and time-consuming (or even impossible for very large datasets). An alternative approach is to use clustering evaluation measures, which can be either external (supervised) or internal (unsupervised) [20]. The former, as the name suggests, use external information (e.g., classification labels) as the ground truth cluster labels. Well known external evaluation measures are Normalised Mutual Information (NMI) [21], Adjusted Mutual Information (AMI) [22], Adjusted Rand Index (ARI) [23, 24], etc. External information, however, is not typically available in real-world scenarios. In such cases we resort to internal evaluation measures, which are solely based on information intrinsic to the data. Although other internal evaluation measures have been proposed [25, 26], we focus on the most commonly-employed, and successful one [27], which is the silhouette coefficient [1].
@@
Figure 1: Illustration of the elements involved in the computation of the silhouette score $s(x_{i})$ for a given data point $x_{i}$ that belongs to cluster $C_{I}$.
@@
4 Methods
The Silhouette Coefficient provides a score that grades the cluster assignment of a data point. To obtain a single score for all the points $x\in X$, the typical approach is micro-averaging (see §2) that averages all the individual scores. The alternative macro-averaging approach averages first the scores per cluster, and then (macro) average the latter.
4.1 Micro-averaged Silhouette: The typical index
Micro-averaging silhouette at the point level, or sample mean in short, is defined as follows:
@@ 4.1 Micro-averaged Silhouette: The typical index
This is the originally proposed averaging strategy by the study that introduced silhouette [1], the one adopted by scikit-learn, and it is the typical approach employed in literature [11]. However, we show in §5 that it is not effective in the case of imbalanced clusters, which is a very common property of real-world datasets.
4.2 Macro-averaged Silhouette: The overlooked index
When clusters are perfectly balanced, the sample mean is a reasonable aggregation strategy. The assumption of perfectly balanced clusters, however, cannot be guaranteed in the real world, where clusters are often imbalanced. In such cases, and when small clusters also matter (e.g., diagnostic reports about rare medical diseases), we argue that micro-averaging is not effective while macro-averaging is robust. This issue is known in fields such as supervised learning [5], but it has not been studied yet for clustering.
To compute the macro-Silhouette, a score $S_{c}$ is computed for each cluster $C_{i}$ as follows:
@@ 4.2 Macro-averaged Silhouette: The overlooked index
This score measures how compact and well separated a cluster is given a clustering solution. For $K$ clusters in that solution, we end up with a set of $K$ cluster silhouette values {$S_{C_{1}},\ldots,S_{C_{K}}$}. The average of these $K$ scores, defined as the macro-averaged Silhouette, can be used to assess the dataset clustering and is more formally defined as follows:
@@
We note that macro-averaging assumes equal weight between the clusters, but other approaches also apply. The weighted average, for example, where weights reflect the support (i.e., the number of points per cluster, normalised), is closer to micro-averaging in nature. Furthermore, other statistics could be applied, such as the max (or the min), capturing the most (least) compact and well (bad) separated cluster.
4.3 Per-cluster sampling: Efficient and robust macro-Silhouette
@@
The computation of the silhouette coefficient for all the $N$ points in a dataset requires the computation of a pairwise distance matrix at the cost of $\mathcal{O}(N^{2})$ operations. This is demanding in terms of computational and space complexity and, hence, not scalable for large datasets [28]. The typical approach to tackle this problem is to compute the silhouette score using a uniformly selected subsample of the dataset.
In a cluster-imbalanced problem, the typical (uniform across data points) sampling may favour the major cluster and may even disregard completely one of the minor clusters. We argue that this practice contradicts the nature of macro-averaging, which assumes that clusters are equally weighted when averaged. To solve this problem, when macro-averaging is aimed, we propose that sampling takes place per cluster, following the macro-averaging spirit.
@@
More specific, we create a subsample of size $L$ for computing the macro-averaged silhouette score by uniformly selecting a subset of $L/K$ points from each cluster $C_{i}$, where $i=1,\ldots,K$. In this way, we ensure that all the clusters contribute a sufficient number of data points to the subsample, preserving the robustness of macro-Silhouette.
5 Experiments
Figure 2: Synthetic dataset, shown on the left, with four equibalanced clusters. The same space is shown on the right, but the relatively distant cluster B now comprises 5,000 points more, yielding a heavy cluster-imbalance. Silhouette is reported per dataset per aggregation strategy. Micro-averaging increases unreasonably by a large margin.
5.1 Analysis on synthetic data
We created a synthetic dataset consisting of four Gaussian clusters, each with 100 points and a variance of 0.1, shown on the left of Figure 2. The micro- and macro-averaged Silhouette scores are both 0.623, a score that is far from perfect due to the overlapping clusters C and D. The same space is shown on the right, but the lower distant cluster B is now populated with 5,000 points more, resulting in a heavy cluster imbalance. Specifically, the ratio of the smallest to the largest cluster for this dataset is $r=0.02$.
When moving from the (balanced) space on the left to the (imbalanced) space on the right, micro changes considerably. That is because the points added to the distant cluster B directly influence the point-level average value upwards (i.e., a relative increase of 41%). The overlap of the minor clusters C and D is less important in this case, when compared to the well-separated major cluster B. By contrast, macro-average remains robust to this change, because each cluster is equally weighted in the average, disregarding their size. As is shown in Figure 3, the sensitivity of micro-averaging to imbalance becomes apparent early on and it could continue increasing if we continued adding points.
@@ 5.1 Analysis on synthetic data
Figure 3: Silhouette score of the dataset of Figure 2, micro- and macro-averaged, for a varying number of points added.
Per-cluster sampling
Using the same dataset, we assess the robustness of the proposed per-cluster sampling. This is clearly shown in Figure 4, by comparing the typical uniform and the proposed per-cluster sampling of the macro-averaged Silhouette score. The more the imbalance, as we move to the right of the Figure, the more the fluctuations of the macro-averaged Silhouette score computed on a uniform sample of 100 points. The per-cluster sampling, on the other hand, remains robust. Similar fluctuations are observed on uniform sampling and micro-averaging (in red).
Figure 4: Macro-averaged Silhouette score, computed on uniform and per-cluster samples of 100 points, as the size of cluster B of Figure 2 increases.
Major overlapping cluster
When adding points to a distant cluster, we observe an increase of micro-averaged Silhouette as opposed to its macro-averaged counterpart. The situation is different, however, when the cluster we add points to is close to others. Figure 5 depicts this space, where we observe that micro-averaged Silhouette drops in value in this case. Macro-Silhouette, on the other hand, again remains robust. This is mainly because we add points that overlap with nearby clusters, reducing the overall score instead of increasing it, as was the case when we added points to a distant cluster (Figure 3).
Figure 5: Synthetic dataset, equibalanced as in Figure 2 on the left, but now cluster B (to which we add 5,000 points on the right) is very close to clusters D and A.
Estimating $k$ (the Silhouette method)
Silhouette has been suggested as an alternative to the problematic “elbow” method when estimating the number of clusters [7]. That is, the quality of a clustering solution (e.g., the predictions of KMeans) is evaluated for different numbers of clusters. The number that leads to the highest Silhouette score is chosen as the optimal one. We applied this method on a synthetic imbalanced dataset of four isotropic Gaussian blobs, [^2] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_blobs.html three with 100 points and one with 2,300. By undersampling from the major cluster, we also produce an equibalanced version of this space where clusters comprised 100 points each. By applying KMeans, then, with $k$ ranging from 2 to 10, we measured the micro and macro averaged Silhouette per space per $k$, using uniform (for micro) and per-cluster (for macro) sampling of 100 points. As is shown in Figure 6, macro-averaged Silhouette reaches a maximum (blue star) on the ground truth number of clusters (red line) in the imbalanced dataset. Micro-averaged Silhouette is maximised for a different number of clusters. In the undersampled (balanced) version of this dataset, shown on the right, both strategies reach their maximum on the ground truth number of clusters, i.e., four, where the red vertical line is. This result can be explained by the lack of robustness of uniform-sampled micro-averaged Silhouette (see Figure 4), which is deceiving when estimating the number of clusters. The robust per-cluster-sampled macro-averaged Silhouette, on the other hand, is not affected.
(a) Imbalanced
(b) Balanced
Figure 6: Estimating the optimal number of clusters (shown with star) when using micro (orange) and macro (blue) averaged Silhouette. We use both, an imbalanced dataset of four isotropic Gaussians, and an undersampled (balanced) version of the same space. A red vertical line shows the ground truth number of clusters.
5.2 Application on real-world datasets
We employed eight real-world datasets [29] of various types (numeric, time-series, images), sizes (from 150 to more than 500,000 items), dimensionality, and with a varying cluster imbalance. To estimate the latter, we computed the ratio ($r$) of the size of the smallest to the largest cluster. Table 1 displays these datasets, sorted according to their imbalance, which ranged from high ($r=0.03$) to low ($r=1.00$).
These eight datasets are summarised below:
  • pendigits comprises 10,992 pen-based handwritten digits (from 00 to $9$). Each data item is represented by a $16$-dimensional vector containing pixel coordinates. digits also comprises (1,797) images of handwritten digits, but each item is an image of 8x8 pixels, resulting in $d=64$ features.
  • cover type contains cartographic variables for predicting forest cover types. The dataset includes 581,012 samples and 54 features, such as elevation, aspect, slope, and soil type. The cover types are classified into seven categories and it is a highly imbalanced dataset ($r=0.03$).
  • gas sensor consists of measurements from 16 chemical sensors exposed to different gases over a period of several months. The dataset includes 13,910 samples and 128 features. It is used for studying the drift in sensor responses over time and developing algorithms for sensor calibration.
  • wine contains the results of a chemical analysis of wines grown in the same region in Italy but derived from three different cultivars. The analysis determined the quantities of 13 constituents found in each of the three types of wines.
  • iris comprises the lengths and widths of the sepals and petals of iris flowers.
  • glass contains the chemical compositions of glass samples, which are classified into seven types of glass.
  • tcga is a collection of gene expression profiles obtained from RNA sequencing of various cancer samples. It includes $801$ data instances, clinical information, normalised counts, gene annotations, and $6$ cancer types’ pathways.
  • mice consists of the expression levels of 77 proteins/protein modifications that produced detectable signals in the nuclear fraction of the cortex. It includes $1,080$ data points and $8$ eight classes of mice based on the genotype, behaviour, and treatment characteristics.
Evaluation metrics
We compute micro and macro Silhouette scores, as internal validation measures. We also employ external validation measures that use the ground truth labels to assess the clustering solution, presented for completeness. The normalized mutual information (NMI) score measures the similarity between two clusterings by normalizing the mutual information score [30]. A higher score indicates a better match between the cluster labels and the ground truth labels. The adjusted mutual information (AMI) adjusts MI for chance groupings [30]. It measures the agreement between two clustering assignments and is normalized against the entropy of the labels to yield a score between 0 and 1. The adjusted rand index (ARI) measures the similarity between two data clusterings by adjusting Rand Index (RI) to account for the chance grouping of elements [23]. The score ranges from -1 to 1, with higher values indicating better clustering performance.
Table 1: Real-world datasets of varying dimensionality ($d$), size ($N$), number of clusters ($k$), imbalance ratio of smallest to largest cluster ($r$). The average macro (MaS) and micro (MaS) Silhouette score, sampled uniformly and per-cluster respectively, is reported across three runs (st. error of mean), along with NMI, ARI, AMI. Sorted by $r$.
Dataset Type N d k r MaS MiS NMI ARI AMI
Iris Numeric 150 4 3 1.00 0.46±0.01 0.46±0.01 0.66 0.62 0.66
Digits Image 1797 64 10 0.95 0.11±0.01 0.13±0.01 0.69 0.56 0.69
Pendigits Time-series 10992 16 10 0.92 0.24±0.01 0.23±0.01 0.67 0.53 0.67
Mice Protein Numeric 1080 77 8 0.70 0.13±0.01 0.12±0.01 0.26 0.14 0.25
Wine Numeric 178 13 3 0.68 0.30±0.01 0.29±0.01 0.86 0.88 0.86
Gas Sensor Time-series 13910 128 6 0.55 0.22±0.03 0.27±0.01 0.19 0.07 0.19
Glass Numeric 214 9 6 0.12 0.20±0.01 0.31±0.01 0.32 0.17 0.29
Covertype Numeric 110393 54 7 0.03 0.25±0.01 0.08±0.01 0.13 0.05 0.13
Experimental settings
Missing values in datasets were replaced with the mean value of the respective feature. [^3] https://scikit-learn.org/stable/modules/generated/sklearn.impute.SimpleImputer.html. We standardized all the features per dataset, by removing the mean and scaling to unit variance, [^4] https://scikit-learn.org/stable/modules/generated/sklearn.preprocessing.StandardScaler.html. to avoid numerical instabilities in the computations [31]. We trained KMeans per dataset, using the ground truth number of clusters and selecting initial cluster centroids with KMeans++.
5.2.1 Results with fixed ground truth number of clusters
Table 1 presents the evaluation results of KMeans across datasets, setting $k$ to the ground truth number of clusters of each dataset. The external validation scores (NMI, ARI, AMI) are high for balanced datasets ($r>0.9$), they drop in two out of three mild-imbalanced ones ($0.5\leq r\leq 0.7$), and they are overall low for both highly imbalanced datasets. When computing Silhouette, we employ uniform (for micro) and per-cluster (for macro) sampling of 100 points per dataset, repeating three times and reporting the average and the standard error of the mean. We observe that the two indices are very close to each other in the balanced and mild imbalanced zones, with any differences not exceeding the standard error of the mean. By contrast, in the highly imbalanced zone, the two indices are far from each other. The sensitivity of micro-averaging to cluster imbalance (§5.1) can explain the observed difference between the two indices. When MiS is greater (e.g., glass), a major distant cluster may be present while when MaS is greater than MiS (e.g., cover type), a major cluster may be close to other ones.
5.2.2 Results on $k$-estimation
We experimented also with estimating $k$ using the Silhouette method. That is by applying clustering for various $k$ values and assessing the two Silhouette aggregation strategies regarding their ability to yield a maximum score for the ground truth number of clusters. The ground truth number of clusters is shown with $k$ in Table 1 and depicted with a red vertical line in Figure 7. In Figure 7, we see that micro and macro averaging yield the same optimal $k$ in five out of eight datasets, all balanced (iris, pendigits) or mildly imbalanced (mice, wine, gas sensor). In the heavily imbalanced datasets of glass and cover type, as expected, the optimal $k$ differs. Overall, the macro-average yields an optimal $k$ that is the same as the ground truth $k$ in two datasets (digits and wine) while typical micro-averaging yields an optimal $k$ that is the same as the grand truth in one (wine).
(a) iris
(b) digits
(c) pendigits